The diameters of two given circles are as
3 : 4, and the sum of the areas of the circles is equal to the area of a circle whose diameter measures 30 c. Find the diameters of the given circles.

Grade:12th Pass

1 Answers

Raju Garg
13 years ago

area of first circle = 22/7*(3x)2=22/7*9x2=198x2/7

area of second circle=22/7(4x)2=22/7*16x2=352x2/7

total area of circles = 198x2/7+352x2/7 = 550x2/7

area of third circle = 22/7*(30)2=22/7*900=19800/7

total area of circles = area of third circle

550x2/7 = 19800/7

550x2=19800

x2=19800/550=36

x = 6

diameter of circles = 6*3*,6*4*

=18,24